### v1

# Continual Learning with Query-Only Attention

Abstract
Continual learning involves learning from a stream of data without repetition of data points, a scenario that is inherently complex due to distributional shift across tasks. We propose a query-only attention mechanism that discards keys and values, yet preserves the core inductive bias of transformer architectures. In continual learning scenarios, this simplified mechanism significantly mitigates both loss of plasticity and catastrophic forgetting, outperforming baselines such as selective re-initialization. We establish a conceptual link between query-only attention, full transformer attention, and model agnostic meta-learning, framing them as instances of meta-learning. We further provide intuition for why query-based models and attention networks help preserve plasticity in continual settings. Finally, through preliminary Hessian spectrum analysis, we observe that models maintaining higher curvature rank across tasks tend to retain plasticity. Our findings suggest that full attention may not be essential for capturing the benefits of meta-learning in continual learning.
1 Introduction
Continual learning remains a fundamental challenge in deep learning,[25] where models must learn from non-stationary data streams without succumbing to catastrophic forgetting [10] or loss of plasticity [3] . While many existing approaches mitigate forgetting using replay buffers [28], regularization, or selective re-initialization, preserving plasticity—the capacity to adapt to new tasks—remains significantly more elusive.
Recent work has shown that attention network [22] in transformer architectures, originally developed for sequence modeling, exhibit strong performance in continual learning . Notably, attention networks tend to retain plasticity across tasks more effectively than traditional MLPs or CNNs [24]. Motivated by this observation, we ask: Can the core mechanism of transformers—attention—be further simplified while retaining its continual learning benefits?
To this end, we propose a query-only attention mechanism that removes keys and values from the attention layer. Surprisingly, this minimalist design not only retains the benefits of full attention in continual learning but often surpasses it. Our experiments show that query-only attention significantly reduces both catastrophic forgetting and plasticity loss, without relying on task boundaries or selective re initialization.
Moreover, we show a deeper relationship between query-only attention, full attention network and Model-Agnostic Meta-Learning (MAML) [6]. Through empirical comparisons and Hessian-based curvature analysis, we demonstrate that both our model and MAML maintain stable near constant hessian ranks throughout training—a hallmark of models that preserve plasticity and enable rapid adaptation.
While much of continual learning research has focused on catastrophic forgetting, we emphasize that our primary goal is mitigating loss of plasticity—the declining ability of models to acquire new knowledge. In our framework, reduced forgetting emerges as a natural byproduct of improved plasticity, rather than the central objective.
Our key findings are:
  • We provide a conceptual explanation of why the query-based model mitigates both loss of plasticity and catastrophic forgetting by converging to a global solution, rather than task-specific local solutions.
  • We show how the query-based model relates to the full attention mechanism used in transformers.
  • We establish a connection between our approach and meta-learning algorithms [7], in particular showing how it adapts ideas from MAML [6] to continual learning.
  • We analyze Hessian spectra and effective rank [13], demonstrating that if rank decreases across tasks correlates with loss of plasticity.
  • We demonstrate superior performance over multiple baselines in mitigating loss of plasticity.
  • Our results further show that catastrophic forgetting is reduced naturally by the query-based model, although eliminating forgetting is not the primary focus of this work.
2 Related work
Deep neural networks have shown remarkable generalization capabilities on unseen tasks. However, they typically operate under the assumption that the training data is stationary and that all samples are available simultaneously during training. In contrast, online learning assumes that data arrives sequentially in a stream and each data point is observed only once, eliminating the concept of epochs. In such a setting, the model must continuously update its parameters to adapt to incoming data. From the model’s perspective, the data distribution is inherently non-stationary, since not all samples are available at the same time. This leads to two major challenges: catastrophic forgetting and loss of plasticity. Catastrophic forgetting — the degradation of performance on previously learned tasks after training on new ones — has been extensively studied in the literature [15], [9], [18]. A more fundamental and less studied issue is loss of plasticity, the gradual reduction in a model’s ability to learn new tasks altogether [3], [16]
Hence, continual learning faces challenge on two fronts, forward performance / mitigating loss of plasticity and backward performance / mitigating catastrophic forgetting. [2] showed that there exists a tradeoff between the two. We show an alternative view that query_based_cl and related architectures attain a global solution which mitigates both loss of plasticity and catastrophic forgetting simultaneously.
Most papers, analyze one of the two of above challenges, however very few papers tackle both challenges simultaneously. Regularization-based continual learning methods such as Elastic Weight Consolidation (EWC) [10], [27], and Learning without Forgetting [14] were designed primarily to mitigate catastrophic forgetting, i.e., the degradation of previously learned tasks. However, these approaches do not directly address the complementary challenge of loss of plasticity, where the model fails to acquire new tasks altogether. Our focus in this work is specifically on loss of plasticity, where forgetting is a downstream consequence rather than the central phenomenon. For this reason, we compare against baselines that are explicitly targeted at plasticity, including [3] and recent attention-based approaches [24], rather than against EWC-style methods that operate in an orthogonal regime.
One such paper is [5] which uses utility based methodology for handling both catastrophic forgetting and loss of plasticity. Here, the authors are working with unknown task boundaries. However, obtaining the utility is expensive especially in the era of large and very large models. Controlling gradient updates based on weight utility leads to reduced ability to retain old tasks as their number increases. Most importantly, this method is ad-hoc solution for continual learning problem and not a more global solution which can scale to very large number of tasks. Our method contrasts in that it obtains a global solution and thus has no reduction in performance irrespective number of tasks.
The core algorithm we developed is most closely related to the paper [24] which utilizes attention network and replay buffer to handle this challenge. However, attention network has O(n**2) in compute which can be challenging in continual learning setting where data will come rapidly. Here, n is the size of replay buffer. We draw our inspiration from this paper on using replay buffer but make a novel hypothesis that query matrix is all thats needed for continual learning. Our method achieves similar or superior performance compare to full attention and can also do the compute in O(n). The other aspect being [24] does not explain the intriguing phenomenon. Here we carry out a detailed empirical and theoretical analysis on why query only attention works.
Our work reveals deep connection between attention network, in-context learning [4], [26], [1] model agnostic meta learning [6], and metric based meta learning [23], [11], [20] under the paradigm of continual learning.
To understand the mechanism of loss of plasticity in continual learning, we utilize a robust metric of calculating effective rank of hessian matrix as shown in [13], [19]. They show that maintaining plasticity requires the effective rank to remain high rather than decreasing. In our study we pair it with our algorithm of query only attention to justify its efficiency in mitigating loss of plasticity by maintaining a non decreasing effective rank across continual learning setting.
3 Background and preliminaries
Our theoretical analysis builds on several standard components: attention networks, meta-learning (in particular MAML), and $k$-nearest neighbors (KNN). We briefly review each here to fix notation and highlight the connections that will be used in Section 6.
3.1 Attention mechanism
In the standard attention network [22], each query vector $q_{i}$ produces a weighted combination of value vectors $V=\{v_{1},\dots,v_{n}\}$:
$\text{Attn}(q_{i},K,V)\;=\;\sum_{j=1}^{n}\alpha_{ij}\,v_{j},$ (1)
where the weights $\alpha_{ij}$ are obtained from a softmax over query–key dot products:
$\alpha_{ij}\;=\;\frac{\exp\big(q_{i}^{\top}k_{j}/\sqrt{d}\big)}{\sum\limits_{j^{\prime}=1}^{n}\exp\big(q_{i}^{\top}k_{j^{\prime}}/\sqrt{d}\big)}.$ (2)
Here $K=\{k_{1},\dots,k_{n}\}$ are key vectors and $d$ is the feature dimension. This requires computing all pairwise dot products $q_{i}^{\top}k_{j}$, which scales as $\mathcal{O}(n^{2})$ in sequence length $n$. In contrast, our query-only model removes $K$ and $V$, learning task-specific weights $\theta^{\prime}$ directly, while still preserving the interpretation of predictions as weighted combinations over context points.
3.2 Meta-learning
Meta-learning aims to enable rapid adaptation to new tasks from a few support examples [6]. Traditional meta-learning assumes task boundaries and episodic training, and is thus non-continual. Recent work [8] , [21] explores meta-learning for continual learning. Our approach draws inspiration from meta-learning while operating fully online.
Model-agnostic meta-learning
MAML [6] optimizes model parameters through inner- and outer-loop updates:
$\Theta^{\prime}_{i}=\Theta-\alpha\nabla_{\Theta}\mathcal{L}_{\Theta}(\text{support}_{i})$ (3)
$\Theta\leftarrow\Theta-\beta\sum_{i=1}^{m}\nabla_{\Theta}\mathcal{L}_{\Theta^{\prime}_{i}}(\text{query}_{i}).$ (4)
3.3 $k$-Nearest neighbors (KNN)
In $k$-nearest neighbors regression, prediction is based on the $k$ closest points in a support set $\mathcal{S}=\{(x_{i},y_{i})\}_{i=1}^{N}$ under a distance metric $d(\cdot,\cdot)$ (e.g., Euclidean).
Given a query input $x$, let $\mathcal{N}_{k}(x)$ denote the indices of the $k$ nearest neighbors. The kNN prediction is the local average:
$\hat{y}(x)=\frac{1}{k}\sum_{i\in\mathcal{N}_{k}(x)}y_{i}.$ (5)
Thus, kNN regression is a non-parametric, memory-based method where predictions adapt directly from nearby support examples.
3.4 In-context learning
In-context learning has recently been interpreted as an implicit $k$-nearest-neighbors mechanism that emerges in the forward pass of transformers [17]. This perspective provides a key motivation for our work: by modifying attention, we aim to enable continual adaptation without requiring explicit task identifiers.
4 Problem statement
We study the continual learning (CL) setting, where a model receives a continuous stream of data. Each data point is observed once during training, without repetition or epochs.
Formally, the stream is generated from a sequence of tasks $\{T_{1},T_{2},\dots,T_{n}\}$. Each task $T_{i}$ is associated with a (potentially non-stationary) distribution $\mathcal{D}_{i}(x,y)$ over input–label pairs $(x,y)$.
During training:
  • The model does not observe task boundaries or task identities.
  • Samples arrive sequentially, drawn from the evolving distribution.
  • The objective is to update the model online while retaining performance across all tasks.
  • Based on the task at hand, the model may update on a single data point or a batch of data-points.
During inference, when the model is at data point $m$ of task $t$, the goal is to predict the next $n$ points $\{m+1,\dots,m+n\}$ from the same task $t$. The accuracy (or loss) on these future points defines the forward performance. A degradation in forward performance over time is taken as evidence of loss of plasticity.
For backward performance, after the model has been trained up to task $t$, we evaluate it on a previous task tjt-j by sampling $n$ data points from task tjt-j. The degradation in accuracy (or increase in loss) on these past tasks quantifies the extent of catastrophic forgetting.
  • Forward testing (plasticity): The model is evaluated using support examples already present in the replay buffer. No task identifiers are needed, since the buffer itself provides sufficient context for adaptation.
  • Backward testing (forgetting): The model is evaluated on a small held-out set of examples from past tasks. Here task identities are assumed to be known, so that performance on specific earlier tasks can be measured.
During testing:
  • Forward testing (plasticity): When the model is at data point $m$ of task $t$, it is evaluated on the next $n$ points $\{m+1,\dots,m+n\}$ from the same task $t$. Task identities are not available in this evaluation, so the model must rely only on the stream itself. Degradation in this metric across tasks indicates loss of plasticity.
  • Backward testing (forgetting): After the model has been trained up to task $t$, it is evaluated on a previous task tjt-j by sampling $n$ data points from $\mathcal{D}_{t-j}$. Here we assume access to a small buffer of examples from past tasks, and thus task identifiers are known at evaluation time. Degradation in performance on these past tasks quantifies catastrophic forgetting.
5 Method
5.1 Query only model with replay buffer
Drawing the connection from attention networks, meta-learning, KNN and replay buffer we design an architecture which obtains an optimal solution for continual learning task, leading to mitigation of both loss of plasticity and catastrophic forgetting simultaneously. A sample data point is $d=(x,y)\in\mathcal{D}$. Let, $x\in\mathbb{R}^{a}$ and $y\in\mathbb{R}^{b}$ We define a buffer $\mathcal{B}$ containing $n$ data-points. Every time step we construct a support set $\mathcal{S}$ of size $m$ sampled from $\mathcal{B}$ such that $m<=n$ depending on the problem at hand.
Hence, $S\in\mathbb{R}^{m*(a+b)}$
$S=\begin{bmatrix}x_{s1}&y_{s1}\\ x_{s2}&y_{s2}\\ \vdots&\vdots\\ x_{sm}&y_{sm}\end{bmatrix}=\begin{bmatrix}s_{1}\\ s_{2}\\ \vdots\\ s_{m}\end{bmatrix}$
Let, the data-point on which we want to make prediction be query $q_{i}$, such that $q_{i}=x_{i}$. The neural-net model is a single query-matrix $\mathcal{Q_{\theta}}\in\mathbb{R}^{(2a+b)*b}$, for illustration purpose and can have multiple layers.
Algorithm 1 Query-Only Attention with Replay Buffer
1:Stream of tasks $\{T_{1},T_{2},\ldots\}$; replay buffer $\mathcal{B}$ of size $n$; support size $m$; query-only model $Q_{\theta}$; learning rate $\eta$
2:Updated parameters $\theta$
3:Initialize $\theta$, $\mathcal{B}\leftarrow\emptyset$
4:for each incoming sample $(x_{t},y_{t})$ do
5:  Insert $(x_{t},y_{t})$ into buffer $\mathcal{B}$; evict oldest if $|\mathcal{B}|>n$
6:  Sample support set $\mathcal{S}=\{(x_{j},y_{j})\}_{j=1}^{m}\subset\mathcal{B}$
7:  Compute scores $d_{j}\leftarrow Q_{\theta}(x_{t},x_{j},y_{j})$
8:  Prediction $\hat{y}_{t}\leftarrow\sum_{j=1}^{m}d_{j}\,y_{j}$
9:  Loss $\mathcal{L}_{t}\leftarrow\ell(\hat{y}_{t},y_{t})$
10:  Update $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{t}$
11:end for
We can thus write the predictive model as,
$\hat{y_{t}}(x)={\sum\limits_{x_{i},y_{i}\in B}Q_{\theta}(x_{t},x_{i},y_{i})*y_{i}}$ (6)
[…]
5.2 MAML with replay buffer
Note.
The adaptation of MAML with replay buffer is still work in progress. We include it here to illustrate a potential direction for combining meta-learning with large buffers in continual learning, but do not claim it as a finalized or fully validated algorithm. That said, our preliminary results are promising for one of the experiments, and suggest this variant may provide a complementary approach to attention-based or query-only models.
We adapt MAML to the continual learning setting by introducing a large replay buffer. To minimize interference across tasks, the buffer is evenly partitioned into $t$ sub-buffers, one per sampled task. Without such partitioning, the sampled support and query examples from different tasks would overlap excessively, leading to degraded task separation and unstable meta-updates. Each task provides a small support set ($s$) and query set ($q$), which is sufficient for the MAML inner/outer loops.
An important property of this setup is that the support/query sizes remain much smaller than those required by attention-based or query-only models. This makes MAML particularly well-suited for backward evaluation (catastrophic forgetting), since only a small set of examples per task is needed to adapt. Standard MAML inner-loop adaptation and outer-loop meta-updates are then applied on these partitioned tasks.
Algorithm 2 MAML with Replay Buffer (work in progress)
1:Stream $(x_{t},y_{t})$; replay buffer $\mathcal{B}$ of size $N$; tasks $t$; support $s$; query $q$; learning rates $\alpha,\beta$
2:Initialize $\theta$, $\mathcal{B}\leftarrow\emptyset$
3:for each incoming $(x_{t},y_{t})$ do
4:  Insert $(x_{t},y_{t})$ into $\mathcal{B}$; evict oldest if $|\mathcal{B}|>N$
5:  Partition $\mathcal{B}$ into $t$ equal sub-buffers
6:  for each task $i=1\dots t$ do
7:   Sample $s$ support and $q$ queries from sub-buffer $i$
8:   Inner update on support with step size $\alpha$
9:   Outer update on queries with step size $\beta$
10:  end for
11:end for
6 Theoretical discussion
6.1 Global vs local solution
To understand why our algorithm overcomes loss of plasticity and catastrophic forgetting, we first define weighted $k$-nearest neighbors
$\hat{y}(x)=\frac{\sum\limits_{i\in N_{k}(x)}w(x,x_{i})\,y_{i}}{\sum\limits_{i\in N_{k}(x)}w(x,x_{i})},$ (7)
where $w(x,x_{i})$ is a weight function that decreases as the distance $d(x,x_{i})$ increases. Common choices include:
$w(x,x_{i})$ $=\frac{1}{d(x,x_{i})},$ (8)
$w(x,x_{i})$ $=e^{-\alpha\,d(x,x_{i})}.$ (9)
In Equation 7, predictions depend only on neighboring data points and not on any learnable parameters. Thus, weighted kkNN avoids loss of plasticity (no parameters to get stuck in low-rank regions) and catastrophic forgetting (no parameters to overwrite). Performance is fully determined by the support set and chosen distance metric.
Comparing Equation 7 with Equation 6, the difference lies in how the distance metric is obtained: in kkNN it is fixed manually, while in the query-based model it is learned as $\theta$. Once $\theta$ is learned, predictions depend primarily on the support set, so continual adaptation occurs in-context rather than through constant parameter updates. This makes the learning global in nature and independent of any single task which is different from vanilla backpropagation or continual backpropagation algorithm where the model updates its parameters continuously local for each task.
6.2 Model agnostic meta-learning for continual learning
We can rewrite equation 6 as,
$\hat{y}_{t}(x)=\sum_{x_{i},y_{i}\in B}\theta^{\prime}_{t,i}\,y_{i},$ (10)
From Equation 10, predictions depend on task-specific parameters $\theta^{\prime}$ generated at inference time. This parallels the task-specific adaptation in MAML’s inner/outer-loop updates (Equations 3, 4). However, unlike query-based models, MAML was originally designed with task IDs and distributions known, an assumption that breaks in continual learning. In our experiments, we find that using a large replay buffer stabilizes MAML training despite this limitation. Even more interestingly, MAML requires a much smaller support set than query-based or attention networks, which could make it a promising direction for mitigating catastrophic forgetting in future continual learning work where memory efficiency is crucial.
6.3 Relationship to attention network
From Equation 1, the prediction is a linear combination of value vectors, which are themselves transformed representations of the input. Comparing this with Equation 10, we see a strong similarity: both aggregate information from a support set using learned weights. The main difference is that in attention (Equation 2), task specific weights are derived from query–key dot products, whereas in the query-only model they come from a learned distance metric $\theta^{\prime}$.
This equivalence suggests that attention networks can also mitigate loss of plasticity, much like the query-only model. However, computing attention weights requires all pairwise query–key dot products, leading to $\mathcal{O}(n^{2})$ complexity for a support set of size $n$. In contrast, our query-only model only compares the current query with the support set, reducing the complexity to $\mathcal{O}(n)$. This efficiency allows us to scale to larger support sets in continual learning, improving performance without incurring the prohibitive cost of full attention.
6.4 Hessian rank analysis
To study plasticity, we analyze the effective rank of the Hessian [13],[19], defined as
$\mathrm{erank}(H)=\exp\!\Big(-\sum_{i=1}^{n}p_{i}\log p_{i}\Big),\quad p_{i}=\frac{\lambda_{i}}{\sum_{j}\lambda_{j}},$ (11)
where $\{\lambda_{i}\}$ are the eigenvalues of the Hessian. A stable effective non-decreasing effective rank indicates models that preserve plasticity across tasks.
7 Experiments
(a) Forward accuracy (permuted MNIST).
(b) Backward accuracy (permuted MNIST).
Figure 1: Permuted MNIST results (mean over 3 seeds; shaded regions indicate $\pm 1$ standard deviation across seeds).
(a) Forward MSE (SCR).
(b) Backward MSE (SCR).
Figure 2: SCR results (mean over 3 seeds; shaded regions indicate $\pm 1$ standard deviation across seeds).
We evaluate forward (plasticity) and backward (forgetting) performance on two benchmarks: Permuted MNIST (abrupt distribution shifts) and Slowly Changing Regression (SCR) (gradual drift). All results are averaged over 3 seeds; shaded regions show $\pm$1 std. Higher accuracy (classification) / lower MSE (regression) is better.
Baselines and support sizes.
We compare our query-based model to vanilla backpropagation (BP), continual backpropagation (CBP), and attention network. On Permuted MNIST, we report two query-based variants: V1 with support size $1000$ and V2 with support size $200$. The transformer uses support size $100$; larger sets are prohibitive due to its $\mathcal{O}(n^{2})$ attention, whereas our query-only method is $\mathcal{O}(n)$. We also include a MAML-style model with a large replay buffer and a small support set of $10$ (trained on $100$ tasks for cost). On SCR, we use a single query-based model with support size $100$, matching the transformer (both $100$) to isolate algorithmic effects under equal memory/compute.
Observations on Permuted MNIST.
Query-based (V1/V2) achieves the highest forward accuracy and the lowest forgetting, and converges faster than the transformer. The transformer reaches comparable final accuracy but is slower due to quadratic cost. BP shows gradual loss of plasticity; CBP improves stability but still lags. MAML converges quickly with strong accuracy despite being trained on only $100$ tasks.
Observations on SCR.
With equal support ($100$) for query-based and transformer, both sustain low MSE in forward testing. BP gradually loses plasticity, and CBP does not learn well on this benchmark. In backward testing, BP/CBP exhibit severe catastrophic forgetting, while query-based and MAML retain performance.
Note that here, including $y_{i}$ in $Q_{\theta}(x_{t},x_{i},y_{i})$ did not significantly improve performance, so we used only $(x_{t},x_{i})$ to reduce computation. In contrast, for Permuted-MNIST the inclusion of labels $y_{i}$ in $Q_{\theta}$ yielded consistently better performance, and was therefore retained.
Result analysis.
Across both benchmarks, query-based and full attention behave similarly, but query-based converges faster and often performs better because it leverages larger support in $\mathcal{O}(n)$ time (Permuted MNIST), and remains competitive even under equal support (SCR). This highlights practical advantages for continual learning where compute and memory are tight.
7.1 Effective rank
Figure 3: Permuted-MNIST. Mean $\pm$ std over 3 seeds. Effective rank co-varies with forward performance; dips align with reduced plasticity.
The effective rank serves as a proxy for plasticity. We measure it at the start of each task using the Hessian of the final layer only, since full-Hessian computation is infeasible. On Permuted-MNIST, vanilla backpropagation shows a steady drop in effective rank, aligning with loss of plasticity. Other models maintain more stable ranks, indicating preserved plasticity. Note that absolute rank values differ across models due to variation in the size of their final layers, which is the only layer we analyze. Due to computational constraints, this analysis was not extended to SCR.
8 Conclusion
We introduced a query-only attention mechanism for continual learning, showing that it mitigates both loss of plasticity and catastrophic forgetting while reducing computational cost compared to full attention. Our analysis connects query-only models to MAML and full attention through the lens of global vs. local solutions, and further relates them to $k$-nearest neighbors. Hessian rank experiments support the role of curvature in sustaining plasticity. While our evaluation is limited to permuted MNIST and regression benchmarks, the results are promising.
A key limitation is the reliance on a support set, which complicates mitigation of catastrophic forgetting. Future work will extend to larger and more diverse benchmarks, more rigorous theoretical analysis and minimize the reliance on explicit support for the task at hand. These findings highlight that meta-learning principles, even in simplified architectures, can provide a path toward scalable continual learning.
Appendix A Technical Appendices and Supplementary Material
A.1 Broader impacts
This work is primarily foundational research in continual learning. The proposed query-only attention mechanism and accompanying analysis aim to improve the understanding of plasticity and forgetting in neural networks. Potential positive impacts include enabling more efficient and adaptive AI systems, which could reduce retraining costs, improve energy efficiency, and support applications such as robotics, healthcare monitoring, and lifelong personal assistants.
At the same time, continual learning technologies can be misused in domains such as surveillance or profiling, where adaptive models might amplify privacy concerns or biases. While the present work is not directly deployable, these risks highlight the need for responsible use and safeguards in future applications.
Overall, this research contributes theoretical and empirical insights into the foundations of continual learning, with the aim of advancing the field in a transparent and beneficial direction.
A.2 Limitations
Our work has a few important limitations. First, the query-only attention model relies on a support set, which can be restrictive for mitigating catastrophic forgetting in practical continual learning scenarios. Second, our evaluation is limited to two benchmarks (Permuted MNIST and Slowly Changing Regression); broader validation on more complex datasets is needed to confirm generality. Third, while we provide theoretical analysis and intuition linking query-only attention to MAML and kkNN, we do not include formal proofs. We acknowledge these as areas for future work, particularly in extending the experimental scope and strengthening the theoretical foundation.
A.3 Permuted MNIST Setup
The MNIST dataset [12] consists of 60,000 training and 10,000 test images of hand-written digits (099), each represented as a $28\times 28$ grayscale image. To adapt this dataset for continual learning, we make the following modifications:
  • Train/test split. For each task, we use the entire 60,000 original training images. These are further divided into 58,000 images for training and 2,000 images for evaluation within the task. The global MNIST test set is not used directly; instead, we re-sample 2,000 held-out examples per task to serve as test data. This ensures consistency across tasks and keeps evaluation lightweight.
  • Downsampling. To reduce computational cost, all images are downsampled from $28\times 28$ to $7\times 7$, giving 49 input features per image.
  • Task generation. Each task corresponds to a new random permutation of the 49 input pixels. The same permutation is applied consistently to all 60,000 images within a task. Labels remain unchanged.
  • Continual stream. The learner observes tasks sequentially. After completing all 58,000 training pairs of a given permutation, the learner encounters remaining 2000 data-points for testing, following which it encounters the next task (with a new permutation). In total, 7,000 such tasks are generated in the continual stream.
This setup forces the continual learner to adapt to a new input representation (permutation) at the start of each task, while retaining performance on past permutations. It is a widely used benchmark for evaluating both plasticity (ability to adapt to new tasks) and stability (ability to avoid catastrophic forgetting).
Table 1: Permuted MNIST configuration: Query-based CL (V1).
Setting Value
Support size ($|B|$) 1000
Replay buffer size 1000
Distance metric Learned $Q_{\theta}$ (query-only)
Optimizer Adam
Learning rate 1e-4
Weight decay 0.0
Batching (batch size =400=400)
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 7000
Steps per task 150
(Input size, Hidden Size, Output Size, Hidden Layers) ( 108, 100, 1, 9 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 6.5 hours/run
Table 2: Permuted MNIST configuration: Query-based CL (V2).
Setting Value
Support size ($|B|$) 200
Replay buffer size 200
Distance metric Learned $Q_{\theta}$ (query-only)
Optimizer Adam
Learning rate 1e-4
Weight decay 0.0
Batching batch size =400=400
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 7000
Steps per task 150
(Input size, Hidden Size, Output Size, Hidden Layers) ( 108, 100, 1, 9 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 4 hours/run
Table 3: Permuted MNIST configuration: MAML.
Setting Value
(Support size, query size, tasks per iteration) ($|B|$) (10, 10, 5)
Replay buffer size 50000
Optimizer Adam
Outer Learning rate 1e-4
Inner Learning rate 1e-2
Weight decay 0.0
Batching batch size =400=400
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 7000
Steps per task 150
(Input size, Hidden Size, Output Size, Hidden Layers) : ( 49, 100, 10, 3 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 0.5 hours (ran only first 100 tasks)
The attention network for continual learning is directly adapted from [24], which contains the comprehensive details.
Table 4: Permuted MNIST configuration: Attention Network baseline.
Setting Value
Support size ($|B|$) 100
Attention Full self-attention ($\mathcal{O}(n^{2})$)
Replay buffer size 100
Optimizer Adam
Learning rate 5e-4
Weight decay 0.0
Batching batch size =400=400
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 7000
Steps per task 150
Architecture 10 layers, 1 head, d_model=59
(Attention Layers, Attention Heads, Dimension) : ( 10, 1, 59 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 10 hours/run
The vanilla backpropagation and continual backpropagation algorithm is from paper [3], which contains more details.
Table 5: Permuted MNIST configuration: Vanilla Backpropagation.
Setting Value
Support size N/A
Replay buffer size 0 (no replay)
Optimizer Adam
Learning rate 1e-4
Weight decay 0.0
Batching Online (batch size =1=1)
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 7000
Steps per task 150
(Input size, Hidden Size, Output Size, Hidden Layers) : ( 49, 200, 10, 3 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 5.0 hours/run
Table 6: Permuted MNIST configuration: Continuous Backpropagation.
Setting Value
Support size N/A
Replay buffer size 0 (no replay)
Optimizer Adam
Learning rate 1e-4
Weight decay 0
Batching Online (batch size =1=1)
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 7000
Steps per task 150
(Input size, Hidden Size, Output Size, Hidden Layers) : ( 49, 200, 10, 3 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 7.0 hours/run
A.4 Slowly changing regression (SCR)
We adopt the Slowly changing regression (SCR) benchmark introduced by Dohare et al. [3], which was designed to study loss of plasticity in continual learning. In this task, regression targets evolve gradually over time according to smoothly drifting functions, creating a non-stationary data stream. We use the same setup and data-generation procedure as in the original paper, ensuring comparability of results. For full details, we refer readers to the experimental protocol in [3].
Table 7: Slowly changing regression task: Query-based CL.
Setting Value
Support size ($|B|$) 100
Replay buffer size 100
Distance metric Learned $Q_{\theta}$ (query-only)
Optimizer Adam
Learning rate 1e-4
Weight decay 0.0
Batching batch size =1=1
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 1000
Steps per task 10000
(Input size, Hidden Size, Output Size, Hidden Layers) : ( 40, 20, 1, 1 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 7 hours/run
Table 8: Slowly changing regression task: MAML.
Setting Value
(Support size, query size, tasks per iteration) ($|B|$) (10, 10, 5)
Replay buffer size 20000
Optimizer Adam
Inner Learning rate 1e-2
Outer Learning rate 1e-4
Weight decay 0.0
Batching batch size =1=1
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 1000
Steps per task 10000
(Input size, Hidden Size, Output Size, Hidden Layers) : ( 20, 40, 1, 1 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 1 hour to run 100 tasks
Table 9: Slowly changing regression task: Attention Network.
Setting Value
Support size ($|B|$) 100
Replay buffer size 100
Distance metric Learned $Q_{\theta}$ (query-only)
Optimizer Adam
Learning rate 1e-4
Weight decay 0.0
Batching batch size =1=1
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 1000
Steps per task 10000
(Attention Layers, Attention Heads, Dimension) : ( 1, 1, 21 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 14 hours/run
Table 10: Slowly changing regression task: Vanilla Backpropagation
Setting Value
Optimizer Adam
Learning rate 1e-4
Weight decay 0.0
Batching batch size =1=1
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 1000
Steps per task 10000
(Input size, Hidden Size, Output Size, Hidden Layers) : ( 20, 40, 1, 1 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 8 hours/run
Table 11: Slowly changing regression task: Continuous Backpropagation
Setting Value
Optimizer Adam
Learning rate 1e-4
Weight decay 0.0
Batching batch size =1=1
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 1000
Steps per task 10000
(Input size, Hidden Size, Output Size, Hidden Layers) : ( 20, 40, 1, 1 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 11 hours/run